The table below represents the results obtained from an experiment. Using graphical method, determine the average rate of appearance of product C, over the time interval of 10 to 20 seconds.

Time (sec)0510152025C (M)00.40.70.91.01.05


1
Expert's answer
2021-04-19T07:47:50-0400

Average rate of Appearance = "\frac{change In Output}{change In Input}"


Average rate ( 10 to 20 sec) = "\frac{1.01-0.70}{20-10}"


Average rate ( 10 to 20 sec) = "\frac{0.31}{10}" = 0.031